the supplement of an angle is equal to one-half of the angle. Find the number of degrees in each angle.
Answer by checkley75(3666)

x+x/2=180
(2x+x)/2=180 cross multiply.
3x=180*2
3x=360
x=360/120
x=120 degrees for the larger angle.
120/2=60 degrees for the smaller angle.
proof
120+60=180
180=180